A large tank initially contains 220 gal of brine where there is dissolved 35lb of salt. Starting


at time t = 0, a brine containing 3lb of dissolved salt per gallon flows into the tank at the rate of


6gal/min is kept uniform by stirring and the well-stirred mixture simultaneously flows out of the


tank at a slower rate of 3gal/min. How much salt is in the tank at the end of 20 min?

1
Expert's answer
2022-02-15T06:20:02-0500

Initial capacity= 220 gal brine and 35lb of salt


Flow in= 6gal/min


Flow out= 3gal/min


After 20 min=?


Amount of gals contained in the tank given the inflow and outflow= 6gal/min-3gal/min

=3gal/min


V(t)= 220+3t


"\frac{dA}{20}"=("\frac{3lbs}{gal}\times\frac{6gal}{min}" )-("\frac{At lbs}{Vt gal}\times\frac{3gal}{min}")


= 20min(3gal/min)


=60gal